### v4

# What halts the growth of galaxies?

11institutetext: Max Planck Institute for Extraterrestrial Physics, Giessenbachstrasse, 85741 Garching, Germany
What halts the growth of galaxies?
The gas reservoir of galaxies can be altered by outflows driven by star-formation and luminous active galactic nuclei. Jets heating the surroundings of host galaxies can also prevent gas cooling and inflows. Spectacular examples for these three mass displacement channels have been observed, but their importance in transforming the galaxy population depends on the occurrence rates of outflow triggers. We aim to investigate the absolute and relative importance of these three channels. In an observation-driven approach, we combine distribution functions and scaling relations to empirically compare average outflow rates across the galaxy total stellar mass spectrum and across cosmic time. This hinges on local outflow studies which should be extended to systematic, large and diverse samples, and we do not yet consider a halo heating effect by radiation-driven outflows. Our results show, independent of simulations, the dominance of star formation-driven outflows in low-mass galaxies. Massive galaxies today are predominately prevented from growing further by jet heating, while at $z=1-3$ all three processes are approximately similarly important. Over the full mass spectrum and cosmic history, outflows driven by the radiation from active galactic nuclei is never the dominant process.
1 Introduction
Star formation (SF) and the growth of super-massive black holes (SMBH) are processes thought to shape the evolution of galaxies. Evidence for this comes from cosmological simulations, where such feedback mechanisms are effective in reproducing e.g., the luminosity function and color distribution of galaxies (e.g. Croton et al., 2006). However, the observational evidence on the role and importance of the various feedback mechanisms is more complex.
Figure 1: Our workflow starts from distribution functions (top row), which are convolved with scaling relations (middle) into rates of mass displacement (bottom row) at a given galaxy stellar mass and redshift. Shown here is the lowest redshift bin.
Observations are most clear in the rare extremes: very massive systems where jets produced near the SMBH heat intergalactic matter (e.g. see reviews of Fabian, 2012; Heckman & Best, 2014), and bulk outflows of gas associated with rapid SMBH growth as an Active Galactic Nucleus (AGN) (e.g., Cicone et al., 2014). In most systems, disentangling the proposed feedback mechanisms and establishing causality is observationally difficult. This is in part due to the different time-scales AGN and SF act on in comparison to the impact of their activity (e.g. Hickox et al., 2014). Nevertheless, multi-variate analyses of local galaxies suggest that the outflow rates are a function of AGN luminosity, SFR and $M_{\star}$ (e.g. Fluetsch et al., 2019). To circumvent peculiarities of individual systems, large surveys of galaxies have been used to establish statistical differences, for example in AGN and non-AGN host galaxies. Understanding the selection effects is crucial however, as are the limitations in characterizing the host galaxy when it is contaminated with AGN light. Nevertheless, the AGN accretion rate distributions as a function of basic host galaxy properties such as stellar mass ($M_{\star}$) and star formation rate (SFR) over cosmic time has been measured (e.g., Aird et al., 2012; Bongiorno et al., 2012; Aird et al., 2018; Georgakakis et al., 2017). Even more robustly established is the feedback in galaxy clusters by linking AGN luminosity and cluster heating (see below).
In this work, we combine existing measurements to quantify which mechanisms are effective at carrying gas out of a galaxy. We focus on outflows associated with SF activity, AGN jets and AGN luminosity, and investigate their importance as a function of galaxy mass and cosmic time.
2 Methodology
The goal of this work is to compute the average outflow rates of different feedback mechanisms as a function of stellar mass and redshift.
For SF and AGN-related outflows, we rely on the relation of Fluetsch et al. (2019, F19 hereafter). F19 found that the total (molecular, ionized and neutral hydrogen) outflow mass rate can be accurately predicted given stellar mass, SFR and AGN luminosity (see left middle panel of Figure 1):
$\dot{M}/M_{\odot}=(1.29\times\mathrm{SFR}+0.81\times L_{\mathrm{AGN},43})^{1.13}/M_{\star,11}^{0.37}$ (1)
where $L_{\mathrm{AGN,43}}$ is the bolometric AGN luminosity in units of $10^{43}\mathrm{erg}/s$, $M_{\star,11}$ is the stellar mass in units of $10^{11}M_{\odot}$, SFR is the star formation rate in units of $M_{\odot}/\mathrm{yr}$.
2.1 Outflows from star formation
We obtain (specific) SFR distributions (Whitaker et al., 2014; Tomczak et al., 2016; Smit et al., 2014; Schreiber et al., 2015; Salmon et al., 2015; Salim et al., 2007; McLure et al., 2011; Labbé et al., 2013; Karim et al., 2011; Kajisawa et al., 2010; Bauer et al., 2013; Zwart et al., 2014) from the literature compilation of Behroozi et al. (2019). For each of these SFR distributions measured at a certain stellar mass and redshift (shown in top left panel of Figure 1 for $z=0$), we use the F19 relation with $L_{\mathrm{AGN},43}=0$ to compute the outflow rate due to SFR alone. Integration over the SFR probability distribution gives the average outflow rate at a given stellar mass (bottom left panel of Figure 1). We use the variance across measurements as approximate uncertainties.
Figure 2: Mass dependence of outflow channels. Panels indicate different redshift intervals. SF-driven outflows (black error bars, §2.1), AGN-driven outflows (blue, §2.2) and AGN jet cavities preventing inflows (orange, §2.3) are compared. AGN outflows are only weakly dependent of stellar mass. The horizontal dashed line indicates the inverse Hubble time at that redshift. In the two right-most redshift panels, the AGN jet power has been extrapolated (open orange points).
2.2 Outflows radiatively pushed by active galactic nuclei
To compute the outflows associated with luminous AGN, we require the AGN luminosity as a function of stellar mass. Aird et al. (2018) found a power-law specific accretion rate distribution (SARD), $p(\lambda|M_{\star})\propto L_{\mathrm{AGN}}/M_{\star}$ over a range of stellar masses and redshifts. At low redshifts, these measurements are shown in the top middle panel of Figure 1. They start from a galaxy sample and compute for each galaxy the AGN luminosity (from X-ray emission) and stellar mass (from broad-band photometry). However, because of limited survey area, their sample is not sensitive to constraining the bright end of the luminosity function (near $\lambda\geq 1$). There, the uncertainties are large and follow a Gaussian extrapolation, because they assume a two-component Gaussian mixture model to empirically fit the distribution. Georgakakis et al. (2017) uses also large-area surveys, however their methodology is different. Instead of starting with a galaxy sample, they detect a AGN sample in the X-rays and by SED fitting measure the SARD of host galaxies. They also do not assume a SARD shape, but estimate the space density in redshift and SARD bins. They find a universal SARD shape corresponding to a powerlaw with index $-1$, with an exponential cut-off at approximately twice the Eddington limit. Such a limit is also required to match the SARD with the X-ray luminosity function (Aird et al., 2013). Here we generally use the measurements of Aird et al. (2018), and fully propagate the space density uncertainties. When uncertainties become large (¿0.5dex) at the bright end, we impose a exponential cut-off of Georgakakis et al. (2017) to the upper error bars of Aird et al. (2018). We propagate the uncertainties in the relative normalisations. Our combination, which is very similar to the recent results of Laloux et al. (2024), is shown as the dashed error bars in the top middle panel of Figure 1.
After converting the AGN luminosity to an outflow rate with eq. 1 (setting SFR=0), we integrate over the SARD to obtain total outflow rate at a given stellar mass. The outflow rates are shown in the bottom middle panel of Figure 1.
2.3 Bubbles blown by active galactic nuclei
In the centers of galaxy clusters and massive galaxies, radio-emitting Active Galactic Nuclei are common (e.g., Sabater et al., 2019a). Their radio jets are thought to inject heat into the cluster gas that would otherwise cool quickly, and the created hot bubbles rise slowly by boyancy and slowly disperse their energy (see e.g., Fabian, 2012). Outside gas that would otherwise fall into the cluster, condense onto and thereby grow a galaxy, is stalled at large distances (see e.g., Croton et al., 2006). To compute the mass accretion prevented by this mechanism, we combine several correlations.
We start with the AGN radio luminosity function shape determined by Sadler et al. (2002). Comparison of radio luminosity works find that the shape shows little evolution and is consistent between works and radio wavelengths (e.g. Smolčić et al., 2009; Heckman & Best, 2014; Sabater et al., 2019b). The fraction of AGN radio-emitting more than $4\times 10^{23}\mathrm{W}/\mathrm{Hz}$ is mass dependent (Smolčić et al., 2009, their figure 11). The combination is a mass-dependent radio luminosity function, shown in the top right panel of Figure 1. Next, the radio luminosity of AGN correlates well with the power stored in bubbles (e.g. Bîrzan et al., 2008; Cavagnolo et al., 2010; O’Sullivan et al., 2011). This is shown in the second plot in the right column. Integrating over the power distribution, we estimate the total power injected into bubbles by AGN jets. The bubble power is eventually radiatively cooled away by cooling flows. The accretion rate is (Fabian, 1994):
$\dot{M}=\frac{2}{5}\frac{P_{\mathrm{Bubble}}\times\mu m_{\mathrm{P}}}{kT}$ (2)
Here the bubble power is used as the cluster luminosity to be emitted at plasma temperature $T$. The remaining constants are the Boltzmann constant $k$, the proton mass $m_{\mathrm{P}}$ and the mean molecular mass, $\mu\approx 0.6$ in ionized plasmas (Pratt et al., 2019). As cluster temperature $T$ we use the relatively tight relation between stellar mass and total mass of the system (Vikhlinin et al., 2009), combined with the relation of total system mass to temperature (Gonzalez et al., 2013; Kravtsov et al., 2018). These relations are shown in the left and right axes of the third panel in the right column of Figure 1.
The resulting total mass rate is shown in the bottom right panel of Figure 1. We propagate the systematic uncertainties of the cavity-radio power relations, which vary most between works, and the fit uncertainties in the mass-temperature and $M_{500}-M_{\star}$ relation. Unlike the outflows rates in the previous two sections, eq 2 computes the rate of an inflow that could have happened, were it not prevented by the AGN energy injection.
3 Results
Our main result is presented in Figure 2. The importance of SF-driven outflows (black, §2.1), AGN-driven outflows (blue, §2.2) and AGN jet bubbles preventing inflows (orange, §2.3) are compared for different redshift intervals. Points above the dashed horizontal line indicate that the process is important, as the outflow rate exceeds the current stellar mass divided by Hubble time. The main findings are: (1) In massive systems, AGN jets dominate. (2) In low-mass systems SF-driven outflows dominate. (3) At higher redshifts ($z\sim 2$), all three channels appear to be important.
Figure 3: Outflow rate redshift evolution in three mass bins. The lower mass systems (lower panels) are dominated by SF outflows (black), while in the high-mass systems (top panel) at low redshift, AGN jet bubbles (orange) dominate.
Figure 3 presents the same information as a function of redshift. For massive galaxies (top panel), SF-driven outflows was most important at $z>1$, but at present time AGN jets are the dominant channel. For smaller systems (middle and bottom panel), SF-driven outflows always dominate. This figure should be interpreted with care, as galaxies grow in mass over cosmic time and thus do not remain in a single panel.
4 Discussion and conclusion
We presented a framework to compare key channels influencing the growth of galaxies. For this purpose, we combined state-of-the-art distribution functions and scaling relations of outflows driven by SF or AGN and cavities blown by AGN jets. Our approach is analogous to semi-analytic modelers treatment of gas reservoirs that are filled by gravitational collapse. The gas reservoir can be emptied by outflows and gas accretion prevented by the heating of the galaxy intergalactic medium (see e.g., Croton et al., 2006).
In our work, we have however not relied upon simulations. Instead, we obtained our results by combining observed relations collated from a large body of research undertaken over the last decades. As a prediction from Figure 3, we should see AGN-driven outflows dominate are widespread at $z\sim 1-2$, and at high galaxy stellar masses they dominate over SF-driven outflows. This is what is indeed found in high-resolution integral-field observations (e.g. Genzel et al., 2014).
We acknowledge that several of the used relations have considerable caveats. To name but a few examples: A more accurate determination of the AGN activity in the luminous regime is required to accurately determine the total luminosity density (see large gray error bars in the middle bottom panel of Figure 1). We have extrapolated a local outflow relation out to high redshifts and across a wide mass range. The Fluetsch relation needs to be extended to larger, well-selected samples. As discussed in §4.2 of the recent review by Harrison & Ramos Almeida (2024), the scatter in the relation may be very large due to a large fraction of luminous AGN lacking strong outflows (e.g., Ramos Almeida et al., 2022). However, because the population calculation in a given AGN luminosity and stellar mass bin requires the (arithmetic) mean outflow rate, the change is potentially small, and even if not would further reduce the low total population outflow rate we find, not altering our conclusions. While the relation was established at moderate to high-mass galaxies, low mass galaxies are more frequently satellite galaxies that may undergo additional environmental quenching (see e.g., Peng et al., 2010; Bluck et al., 2020). Such additional quenching mechanisms are not considered here. More work is also needed to understand the validity of applying cluster scaling relations down to the galaxy group regime, and the importance of projection effects in cavity energy relations (Eckert et al., 2021).
We focused on AGN-driven gas ejection because it is a common narrative. Nevertheless, feedback by radiation-driven feedback of AGN may be effective differently. For example, the energy injected may induce a similar halo heating effect as radio jets, as suggested in simulations (e.g., Gabor & Bournaud, 2014; Bower et al., 2017; Nelson et al., 2019) and evidenced in hot halo gas found by the Sunyaev–Zel’dovich effect in quasars (e.g., Ruan et al., 2015; Crichton et al., 2016; Jones et al., 2023). To address this, the presented framework could be extended in the future with a hybrid of the radio jet heating and AGN outflow calculations. Indeed, a similar work to ours was recently conducted by Heckman & Best (2023), which came to similar conclusions. While this work focused on mass outflow rate, Heckman & Best (2023) focused on energy injection, which can address the entropy injected by AGN outflows as an alternative mechanism. Their approach to feedback from massive stars relied on models, while our approach relies on observed outflow rate relations. Both approaches have systematic uncertainties (see e.g., Hardcastle & Croston, 2020). Their jet bubble calculations are similar to ours, but we calculate from there the prevented rate of gas inflow. For AGN radiation-driven outflows, we use the more recent tight relation by Fluetsch et al. (2019), which allows disentangling causation by star formation and AGN. That the two works with different approaches reach the same conclusions is reassuring. The reverse cross-over case, where equatorial radio jets drive molecular outflows (see discussion in §4.2 of Harrison & Ramos Almeida, 2024) could also be considered in the future, but would require further observational studies that establish relations.
In summary, this study finds that transportation of gas out of galaxies driven by luminous AGN is not an effective feedback mechanism. SF-driven outflows are effective in removing gas in the low-mass regime, while AGN jets are effective in the high-mass regime. Unless there are order-of-magnitude under-estimates in our assumptions, outflows driven by AGN radiation that extract gas from galaxies and is thus unavailable for star formation is a sub-dominant galaxy evolution process. Nevertheless, in massive galaxies and especially at $z=1-3$, the outflow rates of all three channels considered in this work are comparable and of almost equal importance.
Acknowledgements.
JB thanks Chris Harrison, Dominique Eckert, Andrea Merloni, Antonis Georgakakis, Mara Salvato and Cristina Ramos Almeida for insightful conversations and feedback on this manuscript. JB thanks the anonymous referee for their helpful comments.